# Pair-density-wave superconductivity and Anderson’s theorem in bilayer nickelates

### v1

@@ C.2.1 Renormalized gap equation and $T_{c}$
$\lim_{\Delta_{0}\rightarrow 0}\frac{1}{\Delta_{0}}\Big[{\widehat{V}}\,F(\bm{k},i\omega_{n})\,{\widehat{V}}^{*}\,{\widehat{\Delta}}^{\dagger}\Big]$ $=$ $-\Big[{\widehat{V}}\,G_{N}G_{N}^{*}\,{\widehat{\Delta}}\,{\widehat{V}}^{*}\,{\widehat{\Delta}}^{\dagger}\Big]$ (S38)
$=$ $-\Big[\big([{\widehat{V}},G_{N}G_{N}^{*}]+G_{N}G_{N}^{*}{\widehat{V}}\big){\widehat{\Delta}}\,{\widehat{V}}^{*}\,{\widehat{\Delta}}^{\dagger}\Big].$
Comparing Eqs. (S37, S38) with the corresponding expressions for the $t_{\perp}=0$ case, we find that finite interlayer hopping generates an additional commutator term, ${\widehat{V}},G_{N}]$ and ${\widehat{V}},G_{N}G_{N}^{*}]$. Therefore, in order for superconductivity to remain robust against disorder (i.e., $\alpha_{\mathrm{dis}}=0$), the condition
$[{\widehat{V}},G_{N}]=0$ (S39)
must be additionally satisfied in addition to the condition $F_{c}=0$ obtained for $t_{\perp}=0$.
Since $G_{N}$ is expressed solely in terms of $\tilde{\tau}_{0}$ and $\tilde{\tau}_{3}$, the above condition is satisfied for disorder channels proportional to $\tilde{\tau}_{0}$ and $\tilde{\tau}_{3}$, which correspond to $\tau_{0}$ and $\tau_{1}$ in the layer basis. Combining this result with the condition in Eq. S31, we conclude that
$\alpha_{\mathrm{dis}}=0,\quad\text{for disorder channels }\tau_{0},\ \tau_{1}.$ (S40)

### v2

@@ C.3.1 Renormalized gap equation and $T_{c}$
$=$ $-\left(\frac{1}{\omega_{n}^{2}+\xi_{+}^{2}}+\frac{1}{\omega_{n}^{2}+\xi_{-}^{2}}\right)\times\Big[1-\frac{1}{4}\mathrm{Tr}({\widehat{F}}_{c}^{\dagger}{\widehat{F}}_{c})\Big].$ (S47)
In Eqs. (S44) and (S46), the terms involving commutators vanish upon taking the trace $\mathrm{Tr}([{\widehat{V}},G_{N}]{\widehat{V}})=\mathrm{Tr}([{\widehat{V}},G_{N}G_{N}^{*}]{\widehat{\Delta}}{\widehat{V}}^{*}{\widehat{\Delta}}^{\dagger})=0$, because $^{\widehat{V}}$ is one of Pauli matrix $\tilde{\tau}_{i}$, and $G_{N}$ and $G_{N}G_{N}^{*}$ are expressed only by $\{\tilde{\tau}_{0},\tilde{\tau}_{3}\}$.
Performing the momentum summation, the effective density of states (DOS) is modified to $\bar{N}(0)=\frac{1}{2}(N_{+}(0)+N_{-}(0))$. Crucially, however, the ratio $u_{n}=\tilde{\omega}_{n}/\tilde{\Delta}_{n}$ remains identical to the $t_{\perp}=0$ case (See Eq. S33),
$u_{n}\equiv\frac{\omega_{n}}{\Delta_{0}}=u_{n}\left[1-\frac{\alpha_{\mathrm{dis}}}{\Delta_{0}}\frac{\Gamma}{|u_{n}|}\right],$ (S48)
with the same depairing coefficient,
$\alpha_{\mathrm{dis}}=\frac{1}{4}\mathrm{Tr}(\tilde{F}_{c}^{\dagger}\tilde{F}_{c}).$ (S49)
C.4.1 Renormalized gap equation and $T_{c}$
We now evaluate the critical temperature in the presence of disorder, $T_{c}$. The linearized gap equation is given by,
$\frac{1}{g}=\frac{T_{c}}{\mathcal{V}}\sum_{\bm{k},\omega_{n}}\left[\frac{\tilde{\Delta}_{0}}{\tilde{\omega}_{n}^{2}+\xi_{+,\bm{k}}^{2}}+\frac{\tilde{\Delta}_{0}}{\tilde{\omega}_{n}^{2}+\xi_{-,\bm{k}}^{2}}\right].$ (S50)
After performing the summation over momentum and Matsubara frequencies, we observe that since the ratio $u_{n}=\tilde{\omega}_{n}/\tilde{\Delta}_{0}$ is derived to be unchanged, the functional form of the $T_{c}$ suppression remains the same with the $t_{\perp}=0$ case,
$\log\left(\frac{T_{c}}{T_{c,0}}\right)=-\frac{\pi}{4}\frac{\alpha_{\mathrm{dis}}}{T_{c,0}}\Gamma.$ (S51)
While the same ratio between $T_{c}/T_{c,0}$ holds, the clean-limit critical temperature $T_{c,0}$ itself is rescaled by the change in the DOS,
$\frac{T_{c,0}(t_{\perp})}{T_{c,0}(t_{\perp}=0)}=\frac{2N(0)}{N_{+}(0)+N_{-}(0)},$ (S52)
where $N_{\pm}(0)$ denotes the DOS of the $\xi_{\pm}$ band at the Fermi level.